# Landauer Principle and Thermodynamics of Computation

### v1

@@ V.2 Trapped Ions
In a quantum regime where the information is encoded in a qubit, one needs to reconstruct the model that is considered in the classical regime to be applicable in the quantum realm. The work Yan et al. (2018) has explored the quantum LP based on trapped ion $Ca^{+}$. The ion is trapped in a linear Paul trap. Here, the LB is evaluated by the analysis of the system-reservoir correlation and the change in entropy during the execution of the erasure process. Trapped ions are considered as one of the perfect platforms for the exploration of quantum thermodynamics with high accuracy An et al. (2015); Huber et al. (2008); Roßnagel et al. (2014). The two internal levels of the system ion are considered as the qubit system and the vibrational modes of the ion as the finite temperature bath. The LP is analyzed in the system by observing the phonon number in the variation of the ion. The authors have confirmed that the LP holds in the quantum regime experimentally.
@@ V.4 Superconducting Technology
An experiment has been performed on a hardware platform namely, superconducting flux logic for analyzing the quantum LP Saira et al. (2020). The double well potential in this process arises due to the Josephson effect and the flux quantization. The bit erasure process is explored in the parametric regime where the approximation of metastable state is valid. It is observed that the process is bounded by LB.
@@ VI Reversible Computation Model and Thermodynamic Interpretation
In this section, we first provide a brief overview of the reversible model of computation Keyes and Landauer (1970); Likharev (1982); Bennett (1982, 1989); Landauer (1961); Feynman (2018); Richard (1986); Zurek (1989b); Raussendorf and Briegel (2001); Bennett and Landauer (1985): Ballistic computer and Brownian computer mainly, followed by their thermodynamic interpretation. We first discuss the Ballistic computer (BLC) proposed by Fredkin and Toffoli Fredkin and Toffoli (1982) and its limitations. Subsequently, we discuss the Brownian computer (BWC) which utilizes thermal fluctuation to perform a computational process.
Before delving into reversible computation and its theromdynamic interpretation it is worthwhile to briefly overview the two key concepts: thermodynamic reversibility and logical reversibility. Thermodynamic reversibility Jarzynski (1997c); Crooks (1998, 1999); Jarzynski (2000); Seifert (2005); Kawai et al. (2007); Crooks (2011); Wolpert et al. (2024) can be defined as follows: A physical process is considered thermodynamically reversible if and only if the time evolution of the probability distribution in the process can be reversed. This reversal should include the time reversal of changes in external parameters, along with the inversion of the signs of both work and heat quantities. On the other hand, logical reversibility Landauer (1996); Bennett (1973, 2003); Sagawa (2014) is defined as follows: A computational process is logically reversible if and only if it is an bijection. In other words, for any output logical state, there is a unique input logical state. Otherwise, it is considered logically irreversible. According to Landauer Landauer (1961), a positive amount of heat emission is inevitable while a logically irreversible process occurs where information is erased or thrown away. These two fundamental concepts are crucial in the analysis of thermodynamic computation processes, underscore the pivotal role of thermodynamics in computational theory. How thermodynamic reversibility influences the heat generation in an logical irreversible process, for example the information erasure is summarized in Table 1.
@@ VI.1 Ballistic Computer
The principle of the ballistic computation model Fredkin and Toffoli (1982) is based on elastic collisions. This model consists of hard spheres that collide between themselves elastically and with fixed reflective barriers. From the input side of the model, as shown in Fig. 7 (starting line here), a huge number of hard spheres (balls) are fired with equal velocity. If ‘1’ is there in the input, a ball is considered in the starting line else no ball for ‘0’. Due to the collision process inside, the ball changes its direction and collides with the other balls. The balls, after a finite number of collisions, reach their finishing point. This signifies the output of the computer. The mirror of this computer is equivalent to the logic gates of our digital computers, and the balls are equivalent to the signals.
@@
Any bijective function is computable in this model, but it will be unable to compute non-conservative (non-bijective) Boolean functions Øgaard (2021). Though it conveys to decrease in the amount of cost in energy, we encounter some drawbacks of this setup. Two main drawbacks of this system are its sensitivity to small perturbative changes, and the second one is related to the collision of the balls. It is quite difficult to make each collision elastic. Furthermore, these collisions result in the thermal randomness in the system.
@@
Quantum effects can stabilize the system from this problem, but it will bring some new instability Benioff (1982). The wave-packet spreading causes instability in the system in the quantum realm. Benioff Benioff (1982) in his work has discussed a quantum version of the BLC, where he has proposed a way to culminate the effect of the noise due to the wave packet spreading by utilizing a time-independent Hamiltonian.
@@ VI.3 Brownian computer: Thermodynamic interpretation
Bennett’s misidentification of BWC as a reversible thermodynamic process can be attributed to the focus on tracking internal energy rather than thermodynamic entropy. Analyzing thermodynamic reversibility solely based on the tracking of internal energy is misleading. The essential condition for verifying whether a process is reversible lies in tracking the total entropy of the system. If the total entropy ($S_{system}+S_{environment}$) remains constant throughout the process, it signifies a thermodynamically reversible process. Bennett and Landauer’s oversight of not tracking the BWC’s total entropy resulted in the misidentification that BWC model is a thermodynamically reversible process.
@@ VII.2.1 Finite Automata: Basic Aspect
First, consider a natural example of an automaton. Imagine a toll gate controlled by a computer. Assume the gate remains closed until the required amount, say $25$ bucks, is paid. Moreover, assume that there are three sets of coins only: $5$, $10$, and $25$ bucks. Now, let us consider a situation where the driver of the vehicle inserts 25 bucks in the sequence (5, 5, 10, 5). The state of the machines evolves as follows:
@@
DFA means the current state and the current symbol uniquely determine the next state. NFA means the same current state and the same current symbol may nondeterministically lead to different next states, but without any probability distribution. PFA means their is a conditional probability distribution over all possible next states given the current state and the current symbol.
FA have different forms like deterministic (DFA: the current state and the current symbol uniquely determine the next state), non-deterministic (NFA: the same current state and the same current symbol may nondeterministically lead to different next states, but without any probability distribution), and probabilistic finite automaton (PFA: there is a conditional probability distribution over all possible next states given the current state and the current symbol). In the stochastic (or probablistic) automaton, the single-valued transition function $\delta$ will be replaced by a conditional probability distribution. One can also observe multiple accept states described in the literature as ‘terminal states’ for the system Lawson (2003). One can even encounter multiple start states in the process.
@@ VII.2.2 FA: Thermodynamic Aspect
Universal computation machine (aka TM, see details below) requires infinite tape, i.e., infinite memory at its disposal, which often makes the model unsuitable for practical applications. On the other hand, finite state machines (FSM), aka FA, albeit non-universal, that use finite resources, are an alternative model as real-world computers have limited resources. The construction of a thermodynamically efficient model Chu and Spinney (2018); Ouldridge and Wolpert (2022); Kardeş and Wolpert (2022) of FSMs has been in the limelight in recent times to estimate the energetic cost of such models.
@@
Following this direction, it has also been addressed Ouldridge and Wolpert (2022) that the cost for the computational characterization of the DFA divides regular language into two classes. One is the invertible local map and the other is the non-invertible local map. In the former case, zero minimal cost is encountered, whereas in the latter case, high cost is encountered.
@@ VII.3.1 Turing Machine: Basic Aspect
At each step of the computation, the state of the TM reads the alphabet in the square where the tape head is placed and subsequently moves on to a new state $q^{\prime}$ (see Fig. 10). It writes a new alphabet ($a^{\prime}$) on the tape, and then moves its tape head either to the left or to the right. This process is repeated until the system attains the accepted state. Mathematically, this map can be expressed as $(q,a)\rightarrow\,(q^{\prime},a^{\prime},d)$, where $d$ denotes the right or left movement of the tape head. For a given TM, the arguments of the transition states are called instantaneous descriptions (IDs) of the TM. One can also encounter TM’s with no halts.
@@ VIII Thermodynamics of Error Correction
In the quantum regime, the initial research was primarily focused on developing quantum codes Steane (1996a, b); Knill and Laflamme (1997); Gottesman (1998a, b); Bennett et al. (1996); Knill and Laflamme (1997); Lieb et al. (1961) that provided a rigorous framework for error correction Bennett et al. (1996); Knill and Laflamme (1997); Calderbank et al. (1998). Now, advanced concepts like fault-tolerant quantum computation Shor (1996); DiVincenzo and Shor (1996); Gottesman (1998a) provides the route-map to the threshold theorem for error correction in the quantum regime Knill et al. (1996); Aharonov and Ben-Or (1997).
@@ VIII.1 Classical Error Correction
In the late 40’s of ${}^{\text{th}}$ century, the seminal work of Shannon Shannon (1948) led to the foundation of this field and was extended by Hamming in his work Hamming (1950). Since then, this field has gained importance for developing better communication protocols. The extent to which error correction (EC) of the missing bits is possible depends on the design of the error-correcting code (ECC). Generally, there exist two types of ECC, they are block code Jafarkhani (2005); Adler et al. (1983); Feltstrom et al. (2009) and convolutional code Dholakia (1994); Forney (1970); Alfarano et al. (2023) as depicted in Fig. 11. Here we focus on a subfield of block code: Linear code. There are other models of error correction codes that are not covered here, interested readers can go through Pless (1978); Hoffman et al. (1991) for further information.
@@ VIII.2 Quantum Error Correction
Classical EC is a well-developed theory based on the demand for better communication systems. One-to-one mapping from classical to quantum error correction (QEC) is not possible as the quantum world has some constraints of its own, like qubits are governed by the no-cloning principle Nielsen and Chuang (2002). As a consequence even the simple repetition code [^3] In the repetition code, the encoding is done as, for example, $00\rightarrow 0$ 000 and $11\rightarrow 1$ 111. It will correct a corrupted state, say $010$ to its majority value, $000$ in this case., which belongs to the linear block code class, does not work in the quantum domain. Additionally, the phenomenon of wavefunction collapse upon measurement makes quantum unique from classical. In the seminal work of Peter Shor Shor (1995) the first QEC protocol was proposed. Shor in his work has demonstrated that quantum information can be encoded by exploiting the idea of entanglement of qubits. Works in this direction Calderbank and Shor (1996); Preskill (1998); Kitaev (1997); Knill et al. (1998); Gottesman (1998a) have demonstrated that one can suppress the error rate in the quantum regime provided the qubits meet some physical conditions.
A general qubit can be represented
@@
where $\alpha$ and $\beta$ represents complex number satisfying $|\alpha|^{2}+|\beta|^{2}=1$. Thus, a qubit has the power to encode information in the infinite number of possible superposition of the computational basis states, which are denoted by $|0\rangle$ and $|1\rangle$. Therefore, the qubits are subjected to an infinite number of errors. However, due to the digitization of the errors using the Pauli operator, the error counts are reduced to two fundamental errors Nielsen and Chuang (2002). One is the $X$-type error, which is the bit flip error (similar to classical), and the other is the $Z_{pau}$-type error, which is the phase error.
@@ VIII.3 Thermodynamic Interpretation
In Vedral (2000), the ``$0"$ and ``$1"$ states are represented by whether a single molecule of an ideal gas is in the LHS or RHS of a box, respectively, like in Bennet’s MD setup. Say, initially the molecule is in the LHS or 00 state. If it expands isothermally, it works $\Delta W=k_{B}T\ln 2$ at the expense of the same amount of free energy stored in it. This increases its entropy by $\Delta S=k_{B}\,\ln 2$.
@@
(5) In the last step, the system $\mathcal{B}_{\alpha}$ is brought back to its initial state by isothermal compression, and it requires work to be done.
In summary, to correct $A_{\alpha}$ or reduce its entropy $B_{\alpha}$’s free energy is wasted just like how a refrigerator works. Now to restore $B_{\alpha}$’s free energy its entropy ($\Delta S=k_{B}\,\ln 2$) has to be dumped into the environment and it needs at least $\Delta W=k_{B}T\ln 2$ amount of work to be done which makes it exactly similar to the LP, thereby consistent with the second law of thermodynamics.
@@ IX.3 Computer as a heat engine
The study of the thermal machine is one of the primary perspectives of thermodynamics in classical Carnot (1872); Martini (1983); Walker et al. (1985a, b); Van Wylen and Sonntag (1985); Reed (1898); Barton (2019); Scovil and Schulz-DuBois (1959); Szilard (1929) and quantum regime Kim et al. (2011); Kosloff (2013); Roßnagel et al. (2016); Martínez et al. (2016); Chattopadhyay and Paul (2019); Uzdin et al. (2015); Chattopadhyay et al. (2021a); Chattopadhyay (2020); Mohan et al. (2024); Santos and Chattopadhyay (2023); Mukhopadhyay et al. (2018); Sur et al. (2024); Das et al. (2019); Naseem et al. (2020); Chattopadhyay et al. (2021b); Pandit et al. (2021); Singh et al. (2020, 2023). In a simple sense, one can convey that the Carnot engine extracts an amount of heat $Q_{hot}$ from the hot reservoir at temperature $T_{hot}$ and transfers an amount of heat $Q_{cold}$ to the sink at temperature $T_{cold}$. The work done to execute this process is $W_{hot\rightarrow cold}=Q_{hot}-Q_{cold}$. Optimal efficiency is observed when $\mathcal{L}_{hot}-\mathcal{L}_{cold}=0$, where $\mathcal{L}_{hot}=-Q_{hot}/T_{hot}$ is the negentropy that is imprinted in the engine. So the loss that occurs in the execution of the process is just throwing away the negentropy of the amount $Q_{cold}/T_{cold}$. The concept of a computer being equivalent to a Carnot cycle has been explored Carnot (1872); Costa de Beauregard (1989); Brillouin (1962); Prigogine and Nicolis (1985) and it is inferred from the results of Brillouin (1962); Costa de Beauregard (1989) that an ideal computer encounters a zero work balance, while the information delivered during a process is $\ln 2\,dl=d\mathcal{L}_{hot}-d\mathcal{L}_{cold}$ (where $dl$ denotes the entropic cost of processing information). In contrast, while the ideal computer’s operation suggests a loss of negentropy, it also emphasizes the potential for harnessing information to perform work, challenging traditional views of thermodynamic limitations. This duality underscores the complex relationship between information and energy in physical systems.

### v2

@@ V.2 Trapped Ions
In a quantum regime where the information is encoded in a qubit, one needs to reconstruct the model that is considered in the classical regime to be applicable in the quantum realm. The work Yan et al. (2018) has explored the quantum LP based on trapped ion $Ca^{+}$. The ion is trapped in a linear Paul trap. Here, the LB is evaluated by the analysis of the system-reservoir correlation and the change in entropy during the execution of the erasure process. Trapped ions are considered as one of the perfect platforms for the exploration of quantum thermodynamics with high accuracy An et al. (2015); Huber et al. (2008); Roßnagel et al. (2014). The two internal levels of the system ion are considered as the qubit system, and the vibrational modes of the ion as the finite temperature bath. The LP is analyzed in the system by observing the phonon number in the variation of the ion. The authors have confirmed that the LP holds in the quantum regime experimentally.
@@ V.4 Superconducting Technology
An experiment has been performed on a hardware platform, namely, superconducting flux logic, for analyzing the quantum LP Saira et al. (2020). The double well potential in this process arises due to the Josephson effect and the flux quantization. The bit erasure process is explored in the parametric regime where the approximation of the metastable state is valid. It is observed that the process is bounded by LB.
@@ VI Reversible Computation Model and Thermodynamic Interpretation
In this section, we first provide a brief overview of the reversible model of computation Keyes and Landauer (1970); Likharev (1982); Bennett (1982, 1989); Landauer (1961); Feynman (2018); Richard (1986); Zurek (1989b); Raussendorf and Briegel (2001); Bennett and Landauer (1985): Ballistic computer and Brownian computer mainly, followed by their thermodynamic interpretation. We first discuss the Ballistic computer (BLC) proposed by Fredkin and Toffoli Fredkin and Toffoli (1982) and its limitations. Subsequently, we discuss the Brownian computer (BWC), which utilizes thermal fluctuation to perform a computational process.
Before delving into reversible computation and its thermodynamic interpretation, it is worthwhile to briefly overview the two key concepts: thermodynamic reversibility and logical reversibility. Thermodynamic reversibility Jarzynski (1997c); Crooks (1998, 1999); Jarzynski (2000); Seifert (2005); Kawai et al. (2007); Crooks (2011); Wolpert et al. (2024); Sagawa (2018) can be defined as follows: A physical process is considered thermodynamically reversible if and only if the time evolution of the probability distribution in the process can be reversed. This reversal should include the time reversal of changes in external parameters, along with the inversion of the signs of both work and heat quantities. On the other hand, logical reversibility Landauer (1996); Bennett (1973, 2003); Sagawa (2014) is defined as follows: A computational process is logically reversible if and only if it is an bijection. In other words, for any output logical state, there is a unique input logical state. Otherwise, it is considered logically irreversible. According to Landauer Landauer (1961), a positive amount of heat emission is inevitable while a logically irreversible process occurs where information is erased or thrown away. These two fundamental concepts are crucial in the analysis of thermodynamic computation processes, underscore the pivotal role of thermodynamics in computational theory. How thermodynamic reversibility influences the heat generation in an logical irreversible process, for example the information erasure is summarized in Table 1.
@@ VI.1 Ballistic Computer
The principle of the ballistic computation model Fredkin and Toffoli (1982) is based on elastic collisions. This model consists of hard spheres that collide between themselves elastically and with fixed reflective barriers. From the input side of the model, as shown in Fig. 7 (starting line here), a huge number of hard spheres (balls) are fired with equal velocity. If ‘1’ is there in the input, a ball is considered in the starting line, else no ball for ‘0’. Due to the collision process inside, the ball changes its direction and collides with the other balls. The balls, after a finite number of collisions, reach their finishing point. This signifies the output of the computer. The mirror of this computer is equivalent to the logic gates of our digital computers, and the balls are equivalent to the signals.
@@
Any bijective function is computable in this model, but it will be unable to compute non-conservative (non-bijective) Boolean functions Øgaard (2021). Though it conveys to decrease in the amount of cost in energy, we encounter some drawbacks of this setup. The two main drawbacks of this system are its sensitivity to small perturbative changes, and the second one is related to the collision of the balls. It is quite difficult to make each collision elastic. Furthermore, these collisions result in the thermal randomness in the system.
@@
Quantum effects can stabilize the system from this problem, but it will bring some new instability Benioff (1982). The wave-packet spreading causes instability in the system in the quantum realm. Benioff Benioff (1982) in his work has discussed a quantum version of the BLC, where he has proposed a way to mitigate the effect of the noise due to the wave packet spreading by utilizing a time-independent Hamiltonian.
@@ VI.3 Brownian Computer: Thermodynamic Interpretation
Bennett’s misidentification of BWC as a reversible thermodynamic process can be attributed to the focus on tracking internal energy rather than thermodynamic entropy. Analyzing thermodynamic reversibility solely based on the tracking of internal energy is misleading. The essential condition for verifying whether a process is reversible lies in tracking the total entropy of the system. If the total entropy ($S_{system}+S_{environment}$) remains constant throughout the process, it signifies a thermodynamically reversible process. Bennett and Landauer’s oversight of not tracking the BWC’s total entropy resulted in the misidentification that the BWC model is a thermodynamically reversible process.
@@ VII.2.1 Finite Automata: Basic Aspect
First, consider a natural example of an automaton. Imagine a toll gate controlled by a computer. Assume the gate remains closed until the required amount, say 252525 bucks, is paid. Moreover, assume that there are three sets of coins only: 555, 101010, and 252525 bucks. Now, let us consider a situation where the driver of the vehicle inserts 25 bucks in the sequence (5, 5, 10, 5). The state of the machines evolves as follows:
@@
DFA means that the current state and the current symbol uniquely determine the next state. NFA means that the same current state and the same current symbol may nondeterministically lead to different next states, but without any probability distribution. PFA means there is a conditional probability distribution over all possible next states given the current state and the current symbol.
FA have different forms like deterministic (DFA: the current state and the current symbol uniquely determine the next state), non-deterministic (NFA: the same current state and the same current symbol may nondeterministically lead to different next states, but without any probability distribution), and probabilistic finite automaton (PFA: there is a conditional probability distribution over all possible next states given the current state and the current symbol). In the stochastic (or probabilistic) automaton, the single-valued transition function $\delta$ will be replaced by a conditional probability distribution. One can also observe multiple accept states described in the literature as ‘terminal states’ for the system Lawson (2003). One can even encounter multiple start states in the process.
@@ VII.2.2 FA: Thermodynamic Aspect
Universal computation machine (aka TM, see details below) requires infinite tape, i.e., infinite memory at its disposal, which often makes the model unsuitable for practical applications. On the other hand, finite state machines (FSM), aka FA, albeit non-universal, that use finite resources, are an alternative model, as real-world computers have limited resources. The construction of a thermodynamically efficient model Chu and Spinney (2018); Ouldridge and Wolpert (2022); Kardeş and Wolpert (2022) of FSMs has been in the limelight in recent times to estimate the energetic cost of such models.
@@
Following this direction, it has also been addressed Ouldridge and Wolpert (2022) that the cost for the computational characterization of the DFA divides regular language into two classes. One is the invertible local map and the other is the non-invertible local map. In the former case, zero minimal cost is encountered, whereas in the latter case, a high cost is encountered.
@@ VII.3.1 Turing Machine: Basic Aspect
At each step of the computation, the state of the TM reads the alphabet in the square where the tape head is placed and subsequently moves on to a new state $q^{\prime}$ (see Fig. 10). It writes a new alphabet ($a^{\prime}$) on the tape, and then moves its tape head either to the left or to the right. This process is repeated until the system attains the accepted state. Mathematically, this map can be expressed as $(q,a)\rightarrow\,(q^{\prime},a^{\prime},d)$, where $d$ denotes the right or left movement of the tape head. For a given TM, the arguments of the transition states are called instantaneous descriptions (IDs) of the TM. One can also encounter TMs with no halts.
@@ VIII Thermodynamics of Error Correction
In the quantum regime, the initial research was primarily focused on developing quantum codes Steane (1996a, b); Knill and Laflamme (1997); Gottesman (1998a, b); Bennett et al. (1996); Knill and Laflamme (1997); Lieb et al. (1961) that provided a rigorous framework for error correction Bennett et al. (1996); Knill and Laflamme (1997); Calderbank et al. (1998). Now, advanced concepts like fault-tolerant quantum computation Shor (1996); DiVincenzo and Shor (1996); Gottesman (1998a) provide the route-map to the threshold theorem for error correction in the quantum regime Knill et al. (1996); Aharonov and Ben-Or (1997).
@@ VIII.1 Classical Error Correction
In the late 40’s of ${}^{\text{th}}$ century, the seminal work of Shannon Shannon (1948) led to the foundation of this field and was extended by Hamming in his work Hamming (1950). Since then, this field has gained importance for developing better communication protocols. The extent to which error correction (EC) of the missing bits is possible depends on the design of the error-correcting code (ECC). Generally, there exist two types of ECC, they are block code Jafarkhani (2005); Adler et al. (1983); Feltstrom et al. (2009) and convolutional code Dholakia (1994); Forney (1970); Alfarano et al. (2023) as depicted in Fig. 11. Here we focus on a subfield of block code: Linear code. There are other models of error correction codes that are not covered here; interested readers can go through Pless (1978); Hoffman et al. (1991) for further information.
@@ VIII.2 Quantum Error Correction
Classical EC is a well-developed theory based on the demand for better communication systems. One-to-one mapping from classical to quantum error correction (QEC) is not possible as the quantum world has some constraints of its own, like qubits are governed by the no-cloning principle Nielsen and Chuang (2002). As a consequence, even the simple repetition code [^4] In the repetition code, the encoding is done as, for example, $\rightarrow 00$ 000 and $\rightarrow 11$ 111. It will correct a corrupted state, say 010010010, to its majority value, 000000000 in this case. This belongs to the linear block code class and does not work in the quantum domain. Additionally, the phenomenon of wavefunction collapse upon measurement makes quantum unique from classical. In the seminal work of Peter Shor Shor (1995) the first QEC protocol was proposed. Shor in his work has demonstrated that quantum information can be encoded by exploiting the idea of entanglement of qubits. Works in this direction Calderbank and Shor (1996); Preskill (1998); Kitaev (1997); Knill et al. (1998); Gottesman (1998a) have demonstrated that one can suppress the error rate in the quantum regime provided the qubits meet some physical conditions. A general state of a qubit is represented as
@@
where $\alpha$ and $\beta$ represents complex number satisfying $|\alpha|^{2}+|\beta|^{2}=1$. Thus, a qubit has the power to encode information in the infinite number of possible superposition of the computational basis states, which are denoted by |0|0$\rangle$| 0 ⟩ and |1|1$\rangle$| 1 ⟩. However, due to the digitization of the errors using the Pauli operator, the error counts are reduced to two fundamental errors Nielsen and Chuang (2002). One is the $X$-type error, which is the bit flip error (similar to classical), and the other is the $Z_{pau}$-type error, which is the phase error.
@@ VIII.3 Thermodynamic Interpretation
In Vedral (2000), the 0``0"‘ ‘ 0 ” and 1``1"‘ ‘ 1 ” states are represented by whether a single molecule of an ideal gas is in the LHS or RHS of a box, respectively, like Bennet’s MD setup. Say, initially the molecule is in the LHS or 0 state. If it expands isothermally, it works $\Delta W=k_{B}T\ln 2$ at the expense of the same amount of free energy stored in it. This increases its entropy by $\Delta S=k_{B}\,\ln 2$.
@@
(5) In the last step, the system $\mathcal{B}_{\alpha}$ is brought back to its initial state by isothermal compression, and it requires work.
In summary, to correct $A_{\alpha}$ or reduce its entropy $B_{\alpha}$’s free energy is wasted just like how a refrigerator works. Now to restore $B_{\alpha}$’s free energy its entropy ($\Delta S=k_{B}\,\ln 2$) has to be dumped into the environment and it needs at least $\Delta W=k_{B}T\ln 2$ amount of work, which makes it exactly similar to the LP, and thereby consistent with the second law of thermodynamics.
@@ IX.3 Computer as a Heat Engine
The study of the thermal machine is one of the primary perspectives of thermodynamics in classical Carnot (1872); Martini (1983); Walker et al. (1985a, b); Van Wylen and Sonntag (1985); Reed (1898); Barton (2019); Scovil and Schulz-DuBois (1959); Szilard (1929) and quantum regime Kim et al. (2011); Kosloff (2013); Roßnagel et al. (2016); Mart´ınez et al. (2016); Chattopadhyay and Paul (2019); Uzdin et al. (2015); Chattopadhyay et al. (2021a); Chattopadhyay (2020); Mohan et al. (2024); Santos and Chattopadhyay (2023); Mukhopadhyay et al. (2018); Sur et al. (2024); Das et al. (2019); Naseem et al. (2020); Chattopadhyay et al. (2021b); Pandit et al. (2021); Singh et al. (2020, 2023); Gelbwaser-Klimovsky et al. (2013, 2015); Misra et al. (2022). In a simple sense, one can convey that the Carnot engine extracts an amount of heat $Q_{hot}$ from the hot reservoir at temperature $T_{hot}$ and transfers an amount of heat $Q_{cold}$ to the sink at temperature $T_{cold}$. The work done to execute this process is $W_{hot\rightarrow cold}=Q_{hot}-Q_{cold}$. Optimal efficiency is observed when $\mathcal{L}_{hot}-\mathcal{L}_{cold}=0$, where $\mathcal{L}_{hot}=-Q_{hot}/T_{hot}$ is the negentropy that is imprinted in the engine. So the loss that occurs in the execution of the process is just throwing away the negentropy of the amount $Q_{cold}/T_{cold}$. The concept of a computer being equivalent to a Carnot cycle has been explored Carnot (1872); Costa de Beauregard (1989); Brillouin (1962); Prigogine and Nicolis (1985) and it is inferred from the results of Brillouin (1962); Costa de Beauregard (1989) that an ideal computer encounters a zero work balance, while the information delivered during a process is $\ln 2\,dl=d\mathcal{L}_{hot}-d\mathcal{L}_{cold}$ (where $dl$ denotes the entropic cost of processing information). In contrast, while the ideal computer’s operation suggests a loss of negentropy, it also emphasizes the potential for harnessing information to perform work, challenging traditional views of thermodynamic limitations. This duality underscores the complex relationship between information and energy in physical systems.